Circular No. 8289

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2000 CQ_{114}

D. C. Stephens and K. S. Noll, Space Telescope Science Institute; and W. Grundy, Lowell Observatory, report the detection of a binary companion to the transneptunian object (TNO) 2000 CQ_{114} (cf. MPEC 2000-J45, MPO 27660), from observations made on 2003 June 6.003–6.031 UT with camera 2 (NIC2) of the Near-Infrared Camera Multi-Object Spectrometer on the Hubble Space Telescope (HST). The TNO was observed in the F110W and F160W filters (yielding approximately J and H bands), with two exposures in each filter. The two components are well resolved in all four images. PSF fitting was used to determine a separation between the two components of 0′′.178 ± 0′′.005 at p.a. 255^{o}.6 ± 6^{o}; this corresponds to a separation of ≈ 2.4 pixels at the NIC2 pixel scale (0′′.075 per pixel). The spacecraft tracked the motion of 2000 CQ_{114} as it moved > 0′′.25, and the relative position of the two components remained the same during that time. As 2000 CQ_{114} was then 45.544 AU from the earth, the projected separation of the objects in the plane of the sky is 5880 ± 200 km. Magnitudes for the components: primary, [F110W] = 22.243 ± 0.020, [F160W] = 21.787 ± 0.015; secondary, [F110W] = 22.849 ± 0.031, [F160W] = 22.086 ± 0.017. Earlier observations of 2000 CQ_{114} with the HST Wide Field Planetary Camera 3 (scale 0′′.1 per pixel) on 2002 Jan. 6.112–6.139 show no conclusive evidence for a binary, though the data are possibly consistent with an unresolved pair at a separation of ≤ 0′′.12; magnitudes for the combined system in these images are: V = 23.86 ± 0.12; R_{c} = 23.16 ± 0.08; I_{c} = 22.53 ± 0.08.

IRAS 05436–0007 C. Aspin, Gemini Observatory; and B. Reipurth, University of Hawaii,

report on observations made at the Gemini 8-m telescope on Mauna Kea. The near-infrared counterpart to IRAS 05436–0007 (cf. IAUC 8284) was as bright as J = 11.1, H = 9.0, and K_{s} = 7.4 on 2004 Feb. 3 UT in observations with the NIRI instrument. This is 3.6, 3.2, and 2.9 magnitudes brighter in these bands, respectively, than the 2MASS values from 1998 Oct. 7. A K-band NIRI spectrum from the same night shows strong CO band-head emission and the Brγ line in emission. An optical spectrum obtained with the GMOS instrument on 2004 Feb. 14 shows a prominent Hα emission line with P-Cyg profile and a red continuum. These characteristics are consistent with IRAS 05436–0007’s currently undergoing an ‘EXor’ eruption (Herbig 1977, Ap.J. 217, 693; Herbig et al. 2001, PASP 113, 1547).

2004 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green